considered the points a (3,4) and b (7,13). if p is a point on line y=x such that pa+pb is minimum, then the the coordinates of p are

Grade:11

1 Answers

Himanshu Kumar
6 years ago
Answer is (193/26,193/26)As pt say P lies on x=y ist coordinates can be (x,x)Now by distance PA=PB ((x-3)^(2) + (x-4)^(2))^(1/2) = ((x-7)^(2) + (x-13)^(2))^(1/2)On solving we get x=193/26Thus pt. is (193/26,193/26)